Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name E.Chudakov
Log entry time 20:06:00 on December 08, 2008
Entry number 252564
keyword=Moller result
Moller results on Dec 08 17:00-19:30
- The beam tune: the quads were changed, then the beam was sent.
It was at Y=1.5mm in 1H01,4A and 4B. The beam DID NOT move after
the quad change! Perhaps 1.5 mm is not much at this energy.
- The energy lock on Hall C, pulsed, was used.
- The raster was ON.
- Run 13009 was taken with the Hall A laser OFF.
Preliminary results: not corrected for
variations along the targets
# |
beam |
runs |
target |
Ptarget% |
Pbeam% |
1 | 0.65 µA |
13001-13004 | 5 | 7.97 |
54.3±0.3±2.0 |
2 | 0.65 µA |
13005-13008 | 4 | 7.44 |
54.6±0.3±2.0 |
3 | 0.50 µA |
13010-13013 | 2 | 8.21 |
53.9±0.1±2.0 |
average 1-3 | 54.4±0.1±2.0
|
The measured polarization IS LOW for some unknown reason!!!
A copy of this log entry has been emailed to: glamazdi@jlab.org
Figure 1

Figure 2
